OpenStudy (anonymous):

WILL GIVE A MEDAL TO WHOEVER HELPS ME can the set of lengths be the side lengths of a right triangle? 8 ft, 15ft, 17ft yes no

OpenStudy (anonymous):

its no

TheSmartOne (thesmartone):

Yes <--- The answer to your question and technically speaking you can only medal one person. But you can fan as many people as you want.

OpenStudy (anonymous):

I had told u

OpenStudy (anonymous):

we had calculated

TheSmartOne (thesmartone):

8^2+15^2=17^2 64+225=289 289=289

TheSmartOne (thesmartone):

It is a right triangle...

TheSmartOne (thesmartone):

3: 4 :5 6: 8 :10 5: 12 :13 8: 15 :17 7: 24 :25 9: 40 :41 11: 60 :61 12: 35 :37 13: 84 :85 15: 112 :113 16: 63 :65 17: 144 :145 19: 180 :181 20: 21 :29 20: 99 :101 21: 220 :221 24: 143 :145 28: 45 :53 28: 195 :197 32: 255 :257 33: 56 :65 36: 77 :85 39: 80 :89 44: 117 :125 48: 55 :73 51: 140 :149 52: 165 :173 57: 176 :185 60: 91 :109 60: 221 :229 65: 72 :97 84: 187 :205 85: 132 :157 88: 105 :137 95: 168 :193 96: 247 :265 104: 153 :185 105: 208 :233 115: 252 :277 119: 120 :169 120: 209 :241 133: 156 :205 140: 171 :221 160: 231 :281 161: 240 :289 204: 253 :325 207: 224 :305

TheSmartOne (thesmartone):

These are all right triangles^^^

OpenStudy (anonymous):

then u@kahred u made a mistake in calculating 17*17

TheSmartOne (thesmartone):

17*17=289

OpenStudy (anonymous):

okay she would love to take help from u then I am gone wheeeeeee

OpenStudy (anonymous):

THANK U I PASSED

TheSmartOne (thesmartone):

no problem